OpenStudy (anonymous):

what value of c makes the polynomial a perfect square x2+9x+c?

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

to be honest i really dont know, i dont even know where to start but i'm going to use your answero see if it's right

OpenStudy (anonymous):

*to

OpenStudy (a_clan):

is it x squared?

OpenStudy (anonymous):

you divide 9 by 2 and square it. The technique is called Completing the Square.

OpenStudy (a_clan):

we want x^2 + 9x +c in the form of (a+b)^2

OpenStudy (anonymous):

\[c = \big(\frac{9}{2}\big)^2\]

OpenStudy (a_clan):

a^2 + 2ab + b^2 = (a+b)^2 similarly, x^2 +2.(9/2)x +(9/2)^2 = (x+9/2)^2 so, c=(9/2)^2

OpenStudy (anonymous):

thanxx :)